Which of the following has maximum density of nucleus:
 
​a) Si ( A = 30, Z = 14)
​b) P ( A = 31, Z = 15)
c) O ( A = 16, Z = 8)
​d) All have same density. ( Why is d the ans?)

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Radius of a nucles varies with mass number , A, asr=r_0A^{1/3}
Hence density of a nucleus=\dfrac{Am_p}{\dfrac{4}{3}\pi r^3}which becomes independent of A.

So the answer is (d).

Hope it helps. Regards.